What is the probability of throwing a 3 or 6 if you role two dice at once?
In this case, you should first calculate the probability of NOT rolling either a 3 or a 6. For one die, this probability is 4/6 = 2/3. For two dice, this probability is 2/3 x 2/3 = 4/9. Therefore, the probability that you DO get a 3 or a 6 is 1 - 4/9 = 5/9.
What is the probability that a sum of 8 on the two dice will occur?
The probability of rolling (2, 6) is (1/6)2=1/36
Add 1/36 for all the other ways to roll 8 - (3,5),(4,4),(5,3),(6,2)
Total probability is 5/36
The answer depends on how many essay question are in the test.
CANOY a family last name & not canopy is what I am wanting to know.
How do you find the answer to 1 out of 40 with the answer being composite number?
1 out of 40 is a fraction whose absolute value is less than 1. The concepts of prime and composite are relevant for whole numbers (integers) greater than 1. There can, therefore, be no sensible answer to the question.
Find the probability of rolling a 10 with two dice?
Assuming that the random variable is the sum of the two numbers rolled, the answer is 3/36 or 1/12.
Why is 3459 not an even number?
Because 3459 is not completely divisible by 2. Or 3459 is not multiple of 2.
The number of combinations for trumpet players * the combinations for guitarists * the combinations for saxophonists
= (4!/(2!(4-2)!))* (12!/(5!(12-5)!))* (7!/(3!(7-3)!))
= (4!/(2!*2!))* (12!/(5!*7!))* (7!/(3!*4!))
= (6)* ((12*11*10*9*8)/(5*4*3*2))* (7*6*5/(3*2))
= 6*792*35
= 166320
Are m and f mutually exclusive?
If by this you mean male and female then no. Although rare individuals with ambiguous characteristics do occur naturally in human and animal populations. Please see the wikipedia article entitled 'intersex' for more information.
Does variance provide more information than standard deviation?
No. Because standard deviation is simply the square root of the variance, their information content is exactly the same.
Prior probability is the probability that is assessed before reference is made to relevant observations.
For complex events, it is possible to calculate the probability of events, but often extremely difficult. In the given example, for an "average" person (that would need some definition to start with) you would need to know the probability of them scoring a basket without the blindfold - this can be found by observing a number of "average" people attempting a number of baskets and seeing how many are successful (the greater the number of observations, the better the accuracy of the [estimation of the] probability. Also, the effect of blindfolding them needs to be found - this is not so easy, but some measure could possibly be made - and then combining this effect and the probability found some estimation of the probability of the required event can be calculated.
Someone has analysed tennis scoring and given the probability of one of the players winning a point (which can be estimated fairly accurately through past observation) has calculated the probability of them winning the match; however, each match (and even a game within a match) can be affected by further factors (eg one player suffering a small injury) which modify the probability of winning a point, but a calculated probability can still be made.
What are the measurements that fall beyond three standard deviations from the mean?
They are observations with a low likelihood of occurrence. They may be called outliers but there is no agreed definition for outliers.
What is the calulated value of the standard deviation of returns for -22.0 -2.0 20.0 35.0 and 50.0?
It is 28.73
What is a group of happenings that does not affect the probability of the occurrence of the others?
independent
What is one important consequence of reducing standard error of the mean?
Estimates of the mean are then more reliable.
The maximum likelihood estimate, possibly.
What is the probabiltiy of landing on a number with 3 factors on a spinner with 8 possible outcomes?
The answer will depend on what the 8 numbers on the spinner are!
If you go back 34 generations you have 17 billion direct ancestors. How is this possible?
It is not possible. There have never been half that number of people alive at any one time.
This calculation is a crude demonstration, among other things, that everyone is related, however distantly, to everyone else.